OpenStudy (lilbuddy13):

Rationalize the numerator sq. root of x minus sq. root of x+h over h sq. root of x. sq. root of x+h

OpenStudy (lilbuddy13):

looks like this \[\sqrt{x}-\sqrt{x+h} \over h \sqrt{x}\sqrt{x+h}\]

OpenStudy (sshayer):

i have already answered.

OpenStudy (lilbuddy13):

okay thanks

OpenStudy (sshayer):

yw